y' + [1-3x^2/x(1-x^2)]y = 1/1-x^2
1
Expert's answer
2020-05-12T18:31:55-0400

Given differential equation is "y'+ \frac{1-3x^2}{x(1-x^2)}y=\frac{1}{1-x^2} (\clubs)" , clearly, it's a first order first and degree

differential equation of the form "y'+P(x)y=Q(x)" ,Thus integrating factor "(I.F)" is given by,


"I.F=e^{\int P(x) dx}"

Hence, In this case I.F will be,


"I.F=e^{\int \frac{1-3x^2}{x(1-x^2)}dx} \hspace{1.5cm} (\star)"

Now, let's calculate the integral first as follows

"\int \frac{1-3x^2}{x(1-x^2)}dx \hspace{1cm} (\star \star)"

Suppose, "f(x)=x(1-x^2)" ,clearly "f'(x)=1-3x^3" ,thus above integral has the form

"\int \frac{f'(x)}{f(x)}dx"

And we already know that "\int \frac{f'(x)}{f(x)}dx=ln(f(x)) +constant"

Hence, "(\star \star)" will be "ln(x(1-x^2))" ,thus from "(\star)"

"I.F=e^{ln(x(1-x^2)} =x(1-x^2)"

"\big(\because \: e^{ln(f(x))}=f(x)\big)" ,Now on multiply I.F in equation "(\clubs)" we get,


"x(1-x^2)y'+(1-3x^2)y=\frac{x(1-x^2)}{1-x^2}\\ 
\implies \frac{d}{dx}(y \cdot x(1-x^2))=x\\
\implies d(y \cdot x(1-x^2))=x dx\\
\implies \int d(y \cdot x(1-x^2))=\int x dx\\
\implies y \cdot x(1-x^2)=\frac{x^2}{2} +c\\
\implies y=\frac{x}{2(1-x^2)}+\frac{c}{x(1-x^2)}"

where "c" is a constant. Hence, we are done.